what is the area of a section of pavement that is 10ft wide and 70 yd long? give your answer in square feet
Answer by stanbon(75887)

what is the area of a section of pavement that is 10ft wide and 70 yd long? give your answer in square feet
----
Area = (10 ft)(3*70 ft) = 30*70 = 2100 sq. ft
=======================
Cheers,
Stan H.